40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Feb 23 07:14:58 2013, yuta, Update, LSC, can't lock Y arm TRYPOY.png
    Reply  Sat Feb 23 14:52:03 2013, Jenne, Update, LSC, ETMYT camera back to normal 
    Reply  Sat Feb 23 16:54:24 2013, Jenne, Update, LSC, can't lock Y arm 
       Reply  Sat Feb 23 17:14:59 2013, yuta, Update, LSC, can't lock Y arm 
Message ID: 8149     Entry time: Sat Feb 23 16:54:24 2013     In reply to: 8143     Reply to this: 8150
Author: Jenne 
Type: Update 
Category: LSC 
Subject: can't lock Y arm 

I'm not sure that Yuta had found the real Yarm flashes last night.  When I came in today, the Yarm would flash.  However I found that the SRM was aligned, and if I misaligned it, the Yarm flashes would disappear.  So somehow the beam getting into the cavity was related to the reflection off of the SRM.

Later, I moved the TTs, leaving ITMY and ETMY alone, after having misaligned SRM (and ITMX) and I found flashes.  This wasn't ideal though, since the beam was much, much too high on IPANG (beam was half falling off the top of the lens, although yaw was pretty good).  That was also when I changed out the ETMYT camera the first time around.  The flashes on the new camera were visible, but much dimmer than with the Watec.

I tried locking the Yarm in this state, but I could never achieve a lock, even momentarily.  It almost seemed like I wasn't sending actuation signal out to the coils, although signal appeared all the way through the chain until the LSC signal summed with the local damping signal.  I also switched the LSC output matrix to try actuating on the ITM, but I was also not able to lock then.  I have switched it back to have Yarm actuate on ETMY.  I could see a nice PDH signal on POY, and nice flashes on TRY, but no lock at all.  The trigger was triggering, but still no catching of the lock.  I'm not really sure what's up.

After playing with a non-locking, poorly aligned Yarm, I started over by recentering the beam on IPPOS and IPANG using the TTs, but have not been able to get flashing in the cavity again.  After much fitzing around, I put the Watec back at ETMYT, in hopes that we can see flashes again at some point, since it's more sensitive than the old Sony.  Still no flashes though.

I have to leave, but Yuta and Manasa are here, and so I'm leaving the IFO in their custody.

ELOG V3.1.3-